# Using negative controls to identify causal effects with invalid instrumental variables

### v4

@@ C.1 Additional estimators
$0=\sum^{n}_{i=1}S_{i}m(\textbf{Z}_{i},\mathbf{C}_{i})\{Y_{i}-\beta(\mathbf{A}_{i},\mathbf{C}_{i};\psi^{\dagger})S_{i}-t(\mathbf{Z}_{i},\mathbf{C}_{i};\nu^{\dagger})-b_{0}(\mathbf{C}_{i};\theta^{\dagger}_{1})-b_{1}(\mathbf{C}_{i};\theta^{\dagger}_{1})S_{i}\}$
@@
×$\times\{Y_{i}-\beta(\mathbf{A}_{i},\mathbf{C}_{i};\psi^{\dagger})S_{i}-t(\mathbf{Z}_{i},\mathbf{C}_{i};\nu^{\dagger})-b_{0}(\mathbf{C}_{i};\theta^{\dagger}_{0})-b_{1}(\mathbf{C}_{i};\theta^{\dagger}_{0})S_{i}\}$
@@ C.2 Simulation set-up and results
We conducted a series of simulation studies to first evaluate the robustness properties of the proposed estimators from Section 3.2. Specifically, we generated covariates $C_{1}$ and $C_{2}$ from a Bernoulli distribution with expectation 0.5 and a standard normal distribution respectively; let $\mathbf{C}=(C_{1},C_{2})$. An unmeasured confounder $U$ was also a random binary variable mean with expectation 0.5. Then the sample selection probability was $f(S=1|\mathbf{C})=expit(-0.5+C_{1}+0.6C_{2}+0.5C_{1}C_{2})$, and $Z$ was generated from a Bernoulli distribution where $f(Z=1|\mathbf{C})=expit(0.25C_{1}-0.25C_{2}+0.5C_{1}C_{2})$. Generating $S$ in this manner led to around 60% of individuals in a simulated data set being members of the target population ($S$=1). The exposure was simulated from a Bernoulli distribution with $f(A=1|Z,S=1,\mathbf{C},U)=expit(1-1.5Z-0.75C_{1}-0.3C_{2}-0.5C_{1}C_{2}+U)$ for individuals in the population $S=1$ and was fixed at zero for all others. Finally, we generated the outcome from the distribution
@@
In our simulations, we considered a suite of estimators of the conditional average treatment effect on the treated in the target population, discussed in the previous subsection. Nuisance parameters were estimated using maximum likelihood, and a sandwich estimator was obtained for each of the estimators given above. For comparison, we also considered a doubly robust $g$-estimator that is valid under a conditional exchangeability assumption (Robins,, 1994), where a linear model was fitted for the outcome, and a logistic model for the exposure; both models were adjusted for $Z$, $C_{1}$, $C_{2}$ and a $C_{1}C_{2}$ interaction term. Furthermore, we also implemented a standard two-stage least squares estimator under the assumption that $Z$ is a valid instrument. For both stages we fit models that were linear in the instrument/exposure, $C_{1}$, $C_{2}$ and a $C_{1}C_{2}$ interaction term.
@@
    $t(Z,\mathbf{C};\nu^{\dagger})$, $b_{0}(\mathbf{C};\theta_{0}^{\dagger})$ and $t(Z,\mathbf{C};\nu^{\dagger})$, were misspecified: $\mathcal{M}_{NEM}\cap\mathcal{M}_{4}$ holds.
@@
such that $Z$ has no association with $Y$ in the reference population, but does has an association in the target population. In experiment 7, we considered settings where the association between $Z$ and $A$ is weak; the exposure was now generated as $f(A=1|Z,S=1,\mathbf{C},U)=expit(1-0.25Z-0.75C_{1}-0.3C_{2}-0.5C_{1}C_{2}+U)$ for those with $S=1$. Since the aim was to isolate potential bias due to failure of the identification assumptions, all models in experiments 6 and 7 were correctly specfied.

### v6

@@ C.1 Additional estimators
$0=\sum^{n}_{i=1}S_{i}m(\textbf{Z}_{i},\mathbf{C}_{i})\{Y_{i}-\beta(\mathbf{A}_{i},\mathbf{C}_{i};\psi^{\dagger})S_{i}-t(\mathbf{Z}_{i},\mathbf{C}_{i};\nu^{\dagger})-b_{0}(\mathbf{C}_{i};\theta^{\dagger}_{0})-b_{1}(\mathbf{C}_{i};\theta^{\dagger}_{1})S_{i}\}$
@@
×$\times\{Y_{i}-\beta(\mathbf{A}_{i},\mathbf{C}_{i};\psi^{\dagger})S_{i}-t(\mathbf{Z}_{i},\mathbf{C}_{i};\nu^{\dagger})-b_{0}(\mathbf{C}_{i};\theta^{\dagger}_{0})-b_{1}(\mathbf{C}_{i};\theta^{\dagger}_{1})S_{i}\}$
@@ C.2 Simulation set-up and results
We conducted a series of simulation studies to first evaluate the robustness properties of the proposed estimators from Section 3.2. Specifically, we generated covariates $C_{1}$ and $C_{2}$ from a Bernoulli distribution with expectation 0.5 and a standard normal distribution respectively; let $\mathbf{C}=(C_{1},C_{2})$. An unmeasured confounder $U$ was also generated as a random binary variable mean with expectation 0.5. Then the sample selection probability was $f(S=1|\mathbf{C})=expit(-0.5+C_{1}+0.6C_{2}+0.5C_{1}C_{2})$, and $Z$ was generated from a Bernoulli distribution where $f(Z=1|\mathbf{C})=expit(0.25C_{1}-0.25C_{2}+0.5C_{1}C_{2})$. Generating $S$ in this manner led to around 60% of individuals in a simulated data set being members of the population of interest ($S$=1). The exposure was simulated from a Bernoulli distribution with $f(A=1|Z,S=1,\mathbf{C},U)=expit(1-1.5Z-0.75C_{1}-0.3C_{2}-0.5C_{1}C_{2}+U)$ for individuals in the population $S=1$ and was fixed at zero for all others. Finally, we generated the outcome from the distribution
@@
In our simulations, we considered a suite of estimators of the conditional average treatment effect on the treated in the population of interest, discussed in the previous subsection. Nuisance parameters were estimated using maximum likelihood, and a sandwich estimator was obtained for each of the estimators given above. For comparison, we also considered a doubly robust $g$-estimator that is valid under a conditional exchangeability assumption (Robins,, 1994), where a linear model was fitted for the outcome, and a logistic model for the exposure; both models were adjusted for $Z$, $C_{1}$, $C_{2}$ and a $C_{1}C_{2}$ interaction term. Furthermore, we also implemented a standard two-stage least squares estimator under the assumption that $Z$ is a valid instrument. For both stages we fit models that were linear in the instrument/exposure, $C_{1}$, $C_{2}$ and a $C_{1}C_{2}$ interaction term.
@@
    $t(Z,\mathbf{C};\nu^{\dagger})$, $b_{0}(\mathbf{C};\theta_{0}^{\dagger})$ and $b_{1}(\mathbf{C};\theta_{1}^{\dagger})$, were misspecified: $\mathcal{M}_{NEM}\cap\mathcal{M}_{4}$ holds.
@@
such that $Z$ has no association with $Y$ in the reference population, but does has an association in the population of interest. In experiment 7, we considered settings where the association between $Z$ and $A$ is weak; the exposure was now generated as $f(A=1|Z,S=1,\mathbf{C},U)=expit(1-0.25Z-0.75C_{1}-0.3C_{2}-0.5C_{1}C_{2}+U)$ for those with $S=1$. Since the aim was to isolate potential bias due to failure of the identification assumptions, all models in experiments 6 and 7 were correctly specfied.